CH/Agen/93/04

Agenda

Charissa Group Meeting

To be held on: Wednesday, 14 July 1993, 9.00 am

To be held at: Oxford

Distribution: WNC,NMC,RAC,BRF,GJG,SJAH,RAH,JK,JSL,ACM, VP,WDMR,GT,RPW,DDW,DLW,JZ, B.Brooks,D.Smith,D.Savory, all by e-mail/server

Agenda

Demonstration of Electronics

1. Minutes of Meeting of 13 May 1993
2. Matters Arising
3. Update on NSC/SERC matters  (DLW)
4. Progress reports on Electronics
(i) amplifiers, channel control  (R.A.Hunt)
(ii) preamplifiers  (B.Brooks)
(iii) bias supplies  (D.Smith)
5. Costing of Electronics — update and summary
6. Division of Finance and Final Assembly/Testing
7. Timescales for Completion
8. Data Acquisition — Preliminary Discussion

CAMAC acquisition VME acquisition Trigger Logic (CAMAC?) Software

9. Any other business
10. Date of next meeting

W.N. Catford (Secretary) 6 July 1993